OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4. 1 -1 -5 -25

OpenStudy (anonymous):

Substitute given options in the equation and see for which value the equation holds true

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

3x + 3 = -2x -2 5x = -5 =>x = -1

OpenStudy (anonymous):

@ganibl can you help me with one more?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

ty! :] Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution

OpenStudy (anonymous):

-3x - 9 = -3x -8 Since x co-efficient is same on both sides (that is -3) This equation doesnt have any solution

OpenStudy (anonymous):

thank you can you help with just a few more?

OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x x = 10 x = 4 x = -14 x = -17

OpenStudy (anonymous):

x + 20/2 = 3x Just multiply both sides with 2 :) x + 20 = 6x 5x = 20 :) Now U get the answer :)

OpenStudy (anonymous):

would it be x + 10? @ganibl

OpenStudy (anonymous):

No its x+20 over 2 right so when u multiply x+20 over 2 with 2 u will be remained with x + 20 only

OpenStudy (anonymous):

which do you think is the best answer?

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

and for this one if I did it right it would be d. correct? Eddie has saved up $45 to purchase a new camera from the local store. The sales tax in his county is 7% of the sticker price. Write an equation and solve it to determine the value of the highest priced camera Eddie can purchase with his $45, including the sales tax. Round your answer to the nearest penny. a. x - 0.07x = 45; $48.39 b. 0.7x = 45; x = $64.29 c.7x = 45; x = $6.43 d. x + 0.07x = 45; x = $42.06

OpenStudy (anonymous):

Perfect :)

OpenStudy (anonymous):

I have one last question :3 A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432

OpenStudy (anonymous):

@ganibl

OpenStudy (anonymous):

Substitute the given options and check for which value the equation holds true and that will be ur answer :)

OpenStudy (anonymous):

288?

OpenStudy (anonymous):

or simple way is x - 0.8x = 72 0.2x = 72 x = 72/0.2

OpenStudy (anonymous):

So it should be x = 360 :)

OpenStudy (anonymous):

alright thank you :3 also I want to know if I did this right I got 4/25 Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

3/4x+5/8 = 4x 3*8/4x + 5 = 4x 24 = 16x^2 + 20x 16x^2 + 20x -24 = 0 4x^2 + 5x - 6 =0

OpenStudy (anonymous):

what would the fraction be?

OpenStudy (anonymous):

24/4x + 5 = 4x 24/16x^2 + 20x 6/4x^2 + 6/5x

OpenStudy (anonymous):

How did u get 4/25?

OpenStudy (anonymous):

(3/4 x + 5) / (8) = 4x Cross multiply: 8 (4x) = (3/4 x + 5) 32x = 3/4 x + 5 32x - 3/4x = 5 128/4 x - 3/4x = 5 125/4 x = 5 125x = 20 x = 20/125 x = 4/25

OpenStudy (anonymous):

Ah right :) I misunderstood the question :) I thought it as 3/4x+5/8 instead of (3/4x+5)/8

OpenStudy (anonymous):

U got it perfect :)

OpenStudy (anonymous):

can you double check this one cause im pretty sure its 288 A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432

OpenStudy (anonymous):

can u show ur work :) I am sure its 360 :)

OpenStudy (anonymous):

x = 0.8x + 72 x - 0.8x = 72 x(1 - 0.8) = 72 0.2x = 72 x = 72/0.2 x = 72/2/10 = 720/2 = 360 :)

OpenStudy (anonymous):

If we substitute 360 in the given equation 360 = 0.8(360) + 72 360 = 288 + 72 = 360 It satisfies :)

OpenStudy (anonymous):

If we substitute 288 lets see how it goes 288 = 0.8(288) + 72 288 = 230.4 + 72 = 302.4 and is not equal to 288 :) So thats not ur answer

OpenStudy (anonymous):

I just wrote it down ur correct :D thank you!

OpenStudy (anonymous):

No Problem :)